# Token Hidden Reward: Steering Exploration-Exploitation in Group Relative Deep Reinforcement Learning

### v3

@@ 6.4 Generalizing THR to other RL objectives and model families
Combining with other RL objectives. We further show that THR can be seamlessly integrated with other group relative RL objectives. For demonstation, we apply THR to the token level variant of group sequence policy optimization (GSPO-token)  (zheng2025group), which optimizes at the sequence level while allowing token level advantage adjustment (details in Appendix A). Fig. 4 shows that THR($p<0$) boosts Pass@K performance across all K with an average improvement $\sim$0.9% to THR($p=0$) and 1.4% to GSPO. See Apx. for detailed results.
Performance on Llama. To further demonstrate the generality of THR across model families, we evaluate it on Llama3.2-3B-Instruct. Unlike Qwen, Llama exhibits weaker mathematical knowledge, limited cognitive behaviors (gandhi2025cognitive), and faces reduced reasoning length during training. Despite this, as shown in Fig. 5, THR still substantially boosts exploration, achieving up to a 7% Pass@K improvement compared to GRPO. Setting $p<0$ amplifies these exploration gains even further. While baselines such as COV-KL and Pass@K-mixed also provide exploration improvements, they consistently underperform relative to THR. Reduced response length, results on exploitation, exploration results on each dataset, and more training details are provided in Section C.3.
@@ Appendix A Additional Preliminary
Group Sequential Policy Optimization. Recently,  zheng2025group introduce group sequence policy optimization (GSPO), a new reinforcement learning algorithm for training large language models. Following the basic principle of importance sampling, GSPO defines importance ratios based on sequence likelihood and performs sequence-level clipping, rewarding, and optimization. The GSPO objective $\mathcal{J}_{\text{GSPO}}(\theta)$ is then defined as:
@@
Additional Details for Qwen2.5-0.5B-Ins: For the 0.5B model, training is conducted on two A6000 GPUs with a batch size of 32, a maximum rollout length of 2500 tokens, a learning rate of $5e^{-7}$, and a mini-batch size of 16—resulting in two iteration updates per training step. For the greedy decoding performance, we report the best accuracy across multiple checkpoints due to significant fluctuations during training. For all other settings, we report the performance at the final checkpoint. In addition to high-THR tokens, we also include those within the top 20% highest-entropy tokens that do not overlap with high-THR (approximate 4.1 % tokens), and keep their advantage unchanged being $\hat{A}_{i,k}$. For formatting, we follow zeng2025simplerl, adopting simple prompts since the model struggles with complex instructions. We use $p=0.2$ and p=0.2p=-0.2 for exploitation and exploration respectively.
@@
Additional Training Details for Llama: For the Llama3.2-3B-Instruct dubey2024llama model, training is carried out on 8 A100 GPUs with a batch size of 256, a maximum rollout length of 3000 tokens, a learning rate of $1\times 10^{-6}$, and a mini-batch size of 16. For greedy decoding, we report the best accuracy across multiple checkpoints due to the substantial fluctuations observed during training, while for all other settings we report results from the final checkpoint. In addition to high-THR tokens, we also include those within the top 20% highest-entropy tokens that do not overlap with high-THR (approximate 3.5 % tokens ), and fix their keep their advantage unchanged being $\hat{A}_{i,k}$. For formatting, we follow zeng2025simplerl, adopting simple prompts since the model struggles with complex instructions.
@@ C.2 Additional Results on GSPO
We further show that THR can be seamlessly integrated with other group relative reinforcement learning objectives. In particular, we apply THR to token level variant of group sequence policy optimization (GSPO-token)  zheng2025group, which optimizes at the sequence level through clipping, rewarding, and optimization while allow token level advantage adjustment (more details in Appendix Appendix A). As reported in Table 7, incorporating THR with $p<0$ yields substantial improvements, boosting Pass@K performance across all K with an average improvement by around 0.9% to THR and 1.4% to GSPO.
@@
Reduced response length. As shown in Fig. 7, the response length of Llama3.2-3B declines rapidly after a few epochs, with the average length dropping from about 1.5K tokens to roughly 650. This reduction may stem from the model’s limited cognitive behaviors gandhi2025cognitive.
@@ D.1 Pass@K as the question level reweighting
chen2025pass; mahdavi2025beyond; walder2025pass develop RLVR objectives that directly target Pass@K optimization. Starting with GRPO’s ancestor, REINFORCE, mahdavi2025beyond; walder2025pass derive reward rescalings by directly optimizing the Pass@K objective. mahdavi2025beyond apply the same rescaling to advantages giving a GRPO version of their approach. These rescalings upweight the gradient contribution of correct responses that constitute “rare successes”—i.e., responses associated with “hard” questions. Crucially, the reweighting is uniform across all tokens and responses for a given question, which we term question-level reweighting. More recently, chen2025pass introduce an appealing alternative to optimizing Pass@K by incorporating the design directly within GRPO’s group structure. Here, we simplify the formulas in chen2025pass and arrive at an explicit formulation of advantage shaping that reveals its question-level nature. Starting from the defined advantages in chen2025pass:
@@ D.2 Relationship between THR and Entropy Regularizer
Under some mild assumptions, optimizing THR plays a similar role as regularizing [^2] The strength and direction are controlled by the value and sign of hyper-parameter $p$ the evolution of the token entropy in a more efficient way. Because, as stated in the main context, THR considers cross-token influence while current analysis on token entropy consider the influence of learning a observing token on itself cui2025entropy. We start from Lemma 1 proposed in cui2025entropy, which is how the Cov-KL regularizer is derived.
Lemma 1 in cui2025entropy: Let the actor policy $\pi_{\theta}$ be a tabular softmax policy, the difference of information entropy given states between two consecutive steps satisfy:
@@
where the operator $\odot$ is the element-wise multiplication of two vectors, $\chi\triangleq\bm{x},\bm{y}_{<k}$ is the context for the prediction of the $k$-th token, and $c$ is a constant for notation conciseness. In the last equation, we reintroduce the input $\chi$ to the notation to remind readers that the entire equation is conditioned on a given context sequence $\chi$. That is an important extension, because most existing works on entropy regularization (e.g., cui2025entropy) only focus on the influence introduced by updating the observing token on itself. In other words, the $\chi$ for $Q$ and $\mathbf{l}$ are identical. The Cov-KL method compared in Table 4 just applies the quantity above to select tokens with high covariances, and then uses the KL penalty to restrict the update of them.
@@
To make the notations concise, we follow the settings in ren2024learning and use $\chi_{o}$ and $\chi_{u}$ to denote the “observing” token and “updating” context, respectively. Then, Equation 14 becomes:
@@
Following deng2025effect, and under the unconstrained features assumption deng2025effect; mixon2022neural, we then represent $\mathbf{l}^{t}(\chi_{o})=\mathbf{W}^{t}\mathbf{h}_{o}$, where $\mathbf{W}\in\mathbb{R}^{V\times d}$ denotes the shared read-out layer and $\mathbf{h}_{o}\in\mathbb{R}^{d\times 1}$ is the feature vector produced by the LLM backbone, conditioned on the context sequence $\chi_{u/o}=\bm{x},\bm{y}_{u/o,<k}$. Note that while $\mathbf{l}^{t}(\chi_{o})$ shares the same $\mathbf{W}^{t}$, the feature vector $\mathbf{h}$ differs across contexts due to variations in input sequences. The difference vector $\mathbf{l}^{t+1}(\chi_{o})-\mathbf{l}^{t}(\chi_{o})\in\mathbb{R}^{V\times 1}$ can then be expressed as:
@@
When $p<0$, tokens with larger THR values receive stronger penalties. Since, in most cases, $\Delta\mathcal{H}^{t}(\chi)$ and THR point in similar directions, this implies that tokens with higher potential entropy change are penalized, closely aligning with the intuition behind Cov-KL. However, as shown in our experiments, THR achieves greater improvements in exploration performance because it explicitly accounts for cross-token influence, rather than relying solely on entropy-based signals on a token’s self-influence, as in COV-KL cui2025entropy.

### v4

@@ 6.4 Generalizing THR to other RL objectives and model families
Combining with other RL objectives. We further show that THR can be seamlessly integrated with other group relative RL objectives. For demonstation, we apply THR to the token level variant of group sequence policy optimization (GSPO-token)  (Zheng et al., 2025), which optimizes at the sequence level while allowing token level advantage adjustment (details in Appendix A). Fig. 4 shows that THR($p<0$) boosts Pass@K performance across all K with an average improvement $\sim$0.9% to THR($p=0$) and 1.4% to GSPO. See Apx. for detailed results.
Performance on Llama. To further demonstrate the generality of THR across model families, we evaluate it on Llama3.2-3B-Instruct. Unlike Qwen, Llama exhibits weaker mathematical knowledge, limited cognitive behaviors (Gandhi et al., 2025), and faces reduced reasoning length during training. Despite this, as shown in Fig. 5, THR still substantially boosts exploration, achieving up to a 7% Pass@K improvement compared to GRPO. Setting $p<0$ amplifies these exploration gains even further. While baselines such as COV-KL and Pass@K-mixed also provide exploration improvements, they consistently underperform relative to THR. Reduced response length, results on exploitation, exploration results on each dataset, and more training details are provided in Section C.3.
@@ 7 Conclusion
Acknowledgments: This work was partially funded by the NSERC Discovery Grant RGPIN-2021-03677, Alliance Grant ALLRP 581098-22, the Natural Science and Engineering Research Council of Canada (NSERC), the Canada CIFAR AI Chairs program, the Canada Research Chair program, an IITP grant funded by MSIT, and the Digital Research Alliance of Canada.
@@ Appendix A Additional Preliminary
Group Sequential Policy Optimization. Recently,  Zheng et al. (2025) introduce group sequence policy optimization (GSPO), a new reinforcement learning algorithm for training large language models. Following the basic principle of importance sampling, GSPO defines importance ratios based on sequence likelihood and performs sequence-level clipping, rewarding, and optimization. The GSPO objective $\mathcal{J}_{\text{GSPO}}(\theta)$ is then defined as:
@@
Additional Details for Qwen2.5-0.5B-Ins: For the 0.5B model, training is conducted on two A6000 GPUs with a batch size of 32, a maximum rollout length of 2500 tokens, a learning rate of $5e^{-7}$, and a mini-batch size of 16—resulting in two iteration updates per training step. For the greedy decoding performance, we report the best accuracy across multiple checkpoints due to significant fluctuations during training. For all other settings, we report the performance at the final checkpoint. In addition to high-THR tokens, we also include those within the top 20% highest-entropy tokens that do not overlap with high-THR (approximate 4.1 % tokens), and keep their advantage unchanged being $\hat{A}_{i,k}$. For formatting, we follow Zeng et al. (2025), adopting simple prompts since the model struggles with complex instructions. We use $p=0.2$ and p=0.2p=-0.2 for exploitation and exploration respectively.
@@
Additional Training Details for Llama: For the Llama3.2-3B-Instruct Dubey et al. (2024) model, training is carried out on 8 A100 GPUs with a batch size of 256, a maximum rollout length of 3000 tokens, a learning rate of $1\times 10^{-6}$, and a mini-batch size of 16. For greedy decoding, we report the best accuracy across multiple checkpoints due to the substantial fluctuations observed during training, while for all other settings we report results from the final checkpoint. In addition to high-THR tokens, we also include those within the top 20% highest-entropy tokens that do not overlap with high-THR (approximate 3.5 % tokens ), and fix their keep their advantage unchanged being $\hat{A}_{i,k}$. For formatting, we follow Zeng et al. (2025), adopting simple prompts since the model struggles with complex instructions.
@@ C.2 Additional Results on GSPO
We further show that THR can be seamlessly integrated with other group relative reinforcement learning objectives. In particular, we apply THR to token level variant of group sequence policy optimization (GSPO-token)  Zheng et al. (2025), which optimizes at the sequence level through clipping, rewarding, and optimization while allow token level advantage adjustment (more details in Appendix Appendix A). As reported in Table 7, incorporating THR with $p<0$ yields substantial improvements, boosting Pass@K performance across all K with an average improvement by around 0.9% to THR and 1.4% to GSPO.
@@
Reduced response length. As shown in Fig. 7, the response length of Llama3.2-3B declines rapidly after a few epochs, with the average length dropping from about 1.5K tokens to roughly 650. This reduction may stem from the model’s limited cognitive behaviors Gandhi et al. (2025).
@@ D.1 Pass@K as the question level reweighting
Chen et al. (2025); Mahdavi et al. (2025); Walder and Karkhanis (2025) develop RLVR objectives that directly target Pass@K optimization. Starting with GRPO’s ancestor, REINFORCE, Mahdavi et al. (2025); Walder and Karkhanis (2025) derive reward rescalings by directly optimizing the Pass@K objective. Mahdavi et al. (2025) apply the same rescaling to advantages giving a GRPO version of their approach. These rescalings upweight the gradient contribution of correct responses that constitute “rare successes”—i.e., responses associated with “hard” questions. Crucially, the reweighting is uniform across all tokens and responses for a given question, which we term question-level reweighting. More recently, Chen et al. (2025) introduce an appealing alternative to optimizing Pass@K by incorporating the design directly within GRPO’s group structure. Here, we simplify the formulas in Chen et al. (2025) and arrive at an explicit formulation of advantage shaping that reveals its question-level nature. Starting from the defined advantages in Chen et al. (2025):
@@ D.2 Relationship between THR and Entropy Regularizer
Under some mild assumptions, optimizing THR plays a similar role as regularizing [^2] The strength and direction are controlled by the value and sign of hyper-parameter $p$ the evolution of the token entropy in a more efficient way. Because, as stated in the main context, THR considers cross-token influence while current analysis on token entropy consider the influence of learning a observing token on itself Cui et al. (2025). We start from Lemma 1 proposed in Cui et al. (2025), which is how the Cov-KL regularizer is derived.
Lemma 1 in Cui et al. (2025): Let the actor policy $\pi_{\theta}$ be a tabular softmax policy, the difference of information entropy given states between two consecutive steps satisfy:
@@
where the operator $\odot$ is the element-wise multiplication of two vectors, $\chi\triangleq\bm{x},\bm{y}_{<k}$ is the context for the prediction of the $k$-th token, and $c$ is a constant for notation conciseness. In the last equation, we reintroduce the input $\chi$ to the notation to remind readers that the entire equation is conditioned on a given context sequence $\chi$. That is an important extension, because most existing works on entropy regularization (e.g., Cui et al. (2025)) only focus on the influence introduced by updating the observing token on itself. In other words, the $\chi$ for $Q$ and $\mathbf{l}$ are identical. The Cov-KL method compared in Table 4 just applies the quantity above to select tokens with high covariances, and then uses the KL penalty to restrict the update of them.
@@
To make the notations concise, we follow the settings in Ren and Sutherland (2025) and use $\chi_{o}$ and $\chi_{u}$ to denote the “observing” token and “updating” context, respectively. Then, Equation 14 becomes:
@@
Following Deng et al. (2025), and under the unconstrained features assumption Deng et al. (2025); Mixon et al. (2022), we then represent $\mathbf{l}^{t}(\chi_{o})=\mathbf{W}^{t}\mathbf{h}_{o}$, where $\mathbf{W}\in\mathbb{R}^{V\times d}$ denotes the shared read-out layer and $\mathbf{h}_{o}\in\mathbb{R}^{d\times 1}$ is the feature vector produced by the LLM backbone, conditioned on the context sequence $\chi_{u/o}=\bm{x},\bm{y}_{u/o,<k}$. Note that while $\mathbf{l}^{t}(\chi_{o})$ shares the same $\mathbf{W}^{t}$, the feature vector $\mathbf{h}$ differs across contexts due to variations in input sequences. The difference vector $\mathbf{l}^{t+1}(\chi_{o})-\mathbf{l}^{t}(\chi_{o})\in\mathbb{R}^{V\times 1}$ can then be expressed as:
@@
When $p<0$, tokens with larger THR values receive stronger penalties. Since, in most cases, $\Delta\mathcal{H}^{t}(\chi)$ and THR point in similar directions, this implies that tokens with higher potential entropy change are penalized, closely aligning with the intuition behind Cov-KL. However, as shown in our experiments, THR achieves greater improvements in exploration performance because it explicitly accounts for cross-token influence, rather than relying solely on entropy-based signals on a token’s self-influence, as in COV-KL Cui et al. (2025).
Appendix E More Studies
E.1 Ablation study on $p$
In this section, we conduct ablation study on $p$.
Ablation Study on $p>0$ for exploitation: For exploitation, we evaluated $p\in\{0,0.05,0.1,0.2\}$. The results in Table 11 show that decreasing $p$ from 0.1 to 0.05 achieves the higher greedy accuracy, outperforming GRPO by 2.8%. This suggests that a milder exploitation strength is more suitable for the Qwen2.5-Math-1.5B model. In contrast, increasing $p$ to 0.2 leads to a slight drop in greedy accuracy compared with $p=0.1$, likely due to excessive exploitation.
Base Model Method Hard Datasets Standard Datasets Total Avg.
AIME25 AIME24 AMC23 Hard Avg. MATH500 Minerva Olympiad Standard Avg.
Qwen2.5-Math-1.5B Base 0.0 3.3 20.0 7.8 39.6 7.7 24.9 24.1 15.9
GRPO 3.3 13.3 57.5 24.7 71.8 29.0 34.1 45.0 34.8
THR 3.3 13.3 55.0 23.9 70.8 32.4 34.1 45.8 34.8
THR (p=0.1p=-0.1) 10.0 13.3 60.0 27.8 70.6 32.0 32.7 45.1 36.4
THR ($p=0.05$) 10.0 13.3 62.5 28.6 71.8 35.7 32.1 46.5 37.6
THR ($p=0.1$) 3.3 13.3 62.5 26.4 71.4 33.1 34.5 46.3 36.3
THR ($p=0.2$) 3.3 13.3 60.0 25.5 71.0 32.7 33.9 45.9 35.7
Ablation Study on $p<0$ for exploration. For exploration, we evaluate $p\in\{0,-0.05,-0.1,-0.2\}$. As shown in Table 12, we observe a consistent exploration trend where all three $p$ can consistently improve the pass@K performance over GRPO, thus reinforcing the conclusion that $p<0$ can enhance exploration.
Method Qwen2.5-math-1.5B Pass@K
1 2 4 8 16 32 64 128 256
AIME 2025
   GRPO 5.9 9.9 15.0 20.5 26.5 33.6 41.5 49.8 56.7
   THR 5.4 9.2 14.1 19.4 25.0 31.7 39.5 48.0 56.7
   THR (p=0.05p=-0.05) 5.9 10.1 15.5 21.2 27.5 34.7 42.0 49.6 60.0
   THR (p=0.1p=-0.1) 6.0 10.1 15.3 20.9 26.8 33.9 41.7 50.0 60.0
   THR (p=0.2p=-0.2) 6.0 10.2 15.6 21.4 28.1 36.2 44.0 49.8 53.3
AIME 2024
   GRPO 11.4 17.7 24.3 30.5 36.7 43.4 50.0 56.0 63.3
   THR 10.6 16.7 23.4 30.2 37.2 44.8 51.9 58.5 63.3
   THR (p=0.05p=-0.05) 11.9 18.2 24.8 31.0 37.5 44.8 52.9 61.6 70.0
   THR (p=0.1p=-0.1) 11.9 18.2 24.9 31.2 37.9 45.3 52.9 61.2 70.0
   THR (p=0.2p=-0.2) 12.2 18.3 24.7 31.2 38.6 47.4 56.6 64.0 70.0
AMC 2023
   GRPO 46.6 59.1 70.0 78.9 85.5 90.2 93.7 96.0 97.5
   THR 44.8 57.8 69.1 78.2 85.1 90.1 93.6 95.9 97.5
   THR (p=0.05p=-0.05) 48.1 60.6 71.2 79.5 85.3 89.8 93.6 97.1 100.0
   THR (p=0.1p=-0.1) 47.9 61.0 72.2 81.1 87.3 91.6 95.1 98.0 100.0
   THR (p=0.2p=-0.2) 50.3 62.4 72.4 80.4 86.4 91.3 95.4 98.3 100.0
Average
   GRPO 21.3 28.9 36.4 43.3 49.6 55.7 61.7 67.3 72.5
   THR 20.3 28.0 35.5 42.6 49.1 55.5 61.7 67.5 72.5
   THR (p=0.05p=-0.05) 22.0 29.6 37.2 43.9 50.1 56.4 62.8 69.4 76.7
   THR (p=0.1p=-0.1) 21.9 29.8 37.5 44.4 50.7 57.3 63.2 69.7 76.7
   THR (p=0.2p=-0.2) 22.8 30.3 37.6 44.3 51.0 58.3 65.3 70.7 74.4
Table 12: Pass@K performance of different $p<0$ for Qwen2.5-math-1.5B.
E.2 Gradient Steps and Convergence
Effective Gradient Steps. We note that a “step” in our setup corresponds to 32 gradient steps. We follow standard GRPO practice and with a prompt batch size of 256 and 8 rollouts per prompt. Then we use a mini-batch size of 64, resulting in 32 gradient steps per step. Therefore, 40 steps corresponds to 1280 gradient steps.
Validation accuracy along Steps. We show the convergence of training by demonstrating the accuracy of validation dataset of MATH (levels 3–5) Hendrycks et al. (2021), as shown in Figure 11, the validation performance continues to improve gradually until around 30-35 steps, after which the increasing is flat, indicating that the model is convergence, thus we use 40 steps for consistency.
Figure 11: Validation accuracy along training of Qwen2.5-Math-1.5B and Qwen2.5-Math-7B
Reward along Steps. For completeness, we include the reward curves in Figure 12. As shown, the reward rises during the early phase and then stabilizes around 0.55 for the 1.5B model and 0.6 for the 7B model, demonstrating that training remains stable throughout. Although dynamic filtering prevents the reported reward from capturing the true correctness of model outputs, it remains a useful proxy for assessing training stability.
Figure 12: Reward (dynamic filtering applied) along training of Qwen2.5-Math-1.5B and Qwen2.5-Math-7B.
E.3 Comparison with clip-high
In this section, we compare against the clip-high baseline Yu et al. (2025) using the recommended clipping value of 0.28. As shown in Table 13, clip-high improves exploration for $K\geq 32$ relative to GRPO. Nevertheless, despite its strength, THR ($p<0$) consistently surpasses clip-high across all $K$, highlighting the effectiveness of THR ($p<0$) in enhancing exploration.
Method Qwen2.5-Math-1.5B Pass@K
1 2 4 8 16 32 64 128 256
AIME 2025
GRPO 5.9 9.9 15.0 20.5 26.5 33.6 41.5 49.8 56.7
Clip-High 5.6 9.6 14.8 20.5 26.6 33.7 41.6 48.4 53.3
THR ($p<0$) 6.0 10.1 15.3 20.9 26.8 33.9 41.7 50.0 60.0
AIME 2024
GRPO 11.4 17.7 24.3 30.5 36.7 43.4 50.0 56.0 63.3
Clip-High 10.8 16.7 23.2 29.8 36.5 44.0 52.1 60.7 70.0
THR ($p<0$) 11.9 18.2 24.9 31.2 37.9 45.3 52.9 61.2 70.0
AMC23
GRPO 46.6 59.1 70.0 78.9 85.5 90.2 93.7 96.0 97.5
Clip-High 47.3 59.9 70.5 78.8 84.9 89.8 93.8 97.3 100.0
THR ($p<0$) 47.9 61.0 72.2 81.1 87.3 91.6 95.1 98.0 100.0
Average
GRPO 21.3 28.9 36.4 43.3 49.6 55.7 61.7 67.3 72.5
Clip-High 21.2 28.7 36.2 43.0 49.3 55.8 62.5 68.8 74.4
THR ($p<0$) 21.9 29.8 37.5 44.4 50.7 57.3 63.2 69.7 76.7
Table 13: Comparing exploration ability with Pass@K@K on Qwen2.5-Math-1.5B across AIME 2024, AIME 2025, and AMC23.
E.4 Study on error-correction behavior.
In this section, we investigate how THR ($p<0$) relates to corrective and self-verifying behaviors. To quantify this, we compute the ratio of reflection-related tokens to the total number of generated tokens. The full list of reflection-related words used for this analysis is provided in Table 15.
Reflection Words
actually although alternating but
correct despite error fix
however incorrect instead mistake
nevertheless nonetheless note realize
realized rethink reconsider still
thinking think though wait
whereas otherwise wrong yet
unless
Table 14: List of reflection-related words.
Qwen2.5-Math-1.5B
Method #Reflection Token / #Token
GRPO 0.34%
THR 0.36%
THR (p=0.1p=-0.1) 0.55%
Table 15: #Reflection Token / #Token ratio for Qwen2.5-Math-1.5B.
We then report the frequency of these tokens in Table 15, which shows that setting $p<0$ increases the presence of reflection tokens. This indicates that THR ($p<0$) can encourage more verification and correction behavior.
E.5 Tokens retained
Model Avg. Ratio Retained
Qwen2.5-Math-1.5B 18%
Qwen2.5-Math-7B 14%
Table 16: Average fraction of tokens retained under the adaptive threshold $\tau$.
The threshold $\tau$ is inherently adaptive, as it is defined as the average influence of a correct response’s token on the likelihoods of all correct responses. We report in Table 16 the average proportion of tokens retained under this threshold for Qwen2.5-Math-1.5B and Qwen2.5-Math-7B. Notably, the 7B model retains fewer high-THR tokens, which is expected: a stronger model possesses more knowledge, is more confident in its answers, and therefore relies on fewer influential tokens.
Appendix F Usage of Large Language Model
In preparing this paper, we made limited use of ChatGPT to support writing and editing. Specifically, LLMs were employed for language polishing, grammar refinement, and rephrasing sentences to improve clarity and readability. Importantly, all technical content, including theoretical analysis, algorithm design, and experimental results, was conceived, implemented, and validated by the authors. LLM outputs were always critically reviewed, verified, and revised before inclusion. No LLM-generated text, figures, or tables were incorporated without careful human oversight.